OpenStudy (anonymous):

Question 6 (Multiple Choice Worth 1 points) Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (anonymous):

f(x)=2x+2 y=2x+2 f^-1(x) is the inverse x=2y+2 y=(4-2)/2 y=1 f^-1(x)=1

OpenStudy (anonymous):

thank you @damoss